OpenStudy (anonymous):

Which of the following shows the correct steps to find the value of 8 to the power of 1 over 3?

OpenStudy (anonymous):

take the cubed root of eight

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

is there anyway you can show me?

OpenStudy (anonymous):

\[\sqrt[3]{8}\]

OpenStudy (anonymous):

so thats all i have to figure that out?

OpenStudy (anonymous):

the reciprocal of a cube root is a number raise to the power of three\[n^3=8\]

OpenStudy (anonymous):

what is that number when you multiply 3 times will give you an 8?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2*2*2=8

OpenStudy (anonymous):

that is what we mean by \[8^{\frac{1}{3}}\]

OpenStudy (anonymous):

then what is the number?

OpenStudy (anonymous):

which number?

OpenStudy (anonymous):

you multiply 3 times?

OpenStudy (anonymous):

that is the cube root of 8...

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

yup that's correct.... \(\ddot\smile\)

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

yup... you solved your own problem...

OpenStudy (anonymous):

thats not one of my answer choices:S

OpenStudy (anonymous):

you are referring to the steps...?

OpenStudy (anonymous):

idk y the computer put 8 but its 27 lol thats what i had to figure out

OpenStudy (anonymous):

what is the original problem that led you to this?

OpenStudy (anonymous):

Which of the following shows the correct steps to find the value of 27 to the power of 1 over 3?

OpenStudy (anonymous):

it was supose to say that but it put 8 instead

OpenStudy (anonymous):

you mean \(\sqrt[3]{27}\)?

OpenStudy (anonymous):

it is the same as \(n^3=27\), what is n?

OpenStudy (anonymous):

the same approach is applied...

OpenStudy (anonymous):

no\[27 \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

its saying find the value for that

OpenStudy (anonymous):

but when you say to the "power", it tells you how many times you have to multiply that number... and it is a fraction... you will just get the root of that number...

OpenStudy (anonymous):

?????? idk math lol im sorry

OpenStudy (anonymous):

it is \(27^{\frac {1}{3}}\)=\(\sqrt[3]{27}\)...

OpenStudy (anonymous):

ohhhhhhhh

OpenStudy (anonymous):

simply 3 x 3 x 3 = 27

OpenStudy (anonymous):

sp all those things mean the same thing

OpenStudy (anonymous):

yup...

OpenStudy (anonymous):

i get it now thank you for your help sorry for being dificult lol

OpenStudy (anonymous):

no problem... :-)